Find the total electrostatic energy of a uniformly charged sphere of total charge q and radius R



1
Expert's answer
2022-01-13T09:22:03-0500

he total electrostatic energy of a uniformly charged sphere of total charge "q" and radius "R":


"U=\dfrac 35\dfrac {q^2}{4\pi\epsilon_0R}."